Which of the following aqueous solutions has the highest boiling point?

A

0.1 M KNO3

B

0.1 M Na3PO4

C

0.1 M BaCl2

D

0.1 M K2SO4

Grade:12th pass

1 Answers

Rituraj Tiwari
3 years ago
Boiling point elevation depends on the solvent, of course, but we're assuming water, and upon the concentration of the solute and how many particles into which the solute breaks. Colligative properties like boiling point elevation, depend on the total number of particles in solution. The solvent used here is water. So elevation in boiling point will be above a boiling point of water for all solutions.
ΔT=i×Kb​×m, wherei= van Hoff's factor,Kb​= boiling point elevation constant,m= molality
HigherΔTwill have a higher boiling point.
For0.1MKNO3​(i)=(1+(n−1)α)=2(αfor saltis1)
Similarly for0.1MBaCl2​(i)=3(αfor salt is1)

For0.1MNa3​PO4​(i)=4(αfor salt is1)

For0.1MK2​SO4​(i)=3(αfor salt is1)
Thus,Na3​PO4​has the highest boiling point.